OpenStudy (anonymous):

This one is quite easy, let us give it a try: Find the no of ordered triplets \( (x, y, z) \), where \( x\),\( y \) and \(z\) are primes and \( x^y + 1 = z \) Please note that your answer should be conclusive.

OpenStudy (anonymous):

Where to start ... (1, 1, 2) takes care of all of the even primes

OpenStudy (anonymous):

1 is a prime number ??!!!

OpenStudy (asnaseer):

no its not - surely?

OpenStudy (anonymous):

Yes, one is neither prime not composite.

jhonyy9 (jhonyy9):

only Goldbach considered the 1 is a prime number

OpenStudy (anonymous):

opps, I only got z prime, sorry ..

OpenStudy (asnaseer):

(2,2,5) is the first solution - but I am trying to construct some sort of generalisation to cover all solutions

jhonyy9 (jhonyy9):

but today 1 not is prime so the first prime is 2

OpenStudy (anonymous):

I don't know what are you talking about jhonny but isn't the conjecture is that every even integer greater than 2 can be expressed as the sum of two primes, how does 1 comes in question ?

jhonyy9 (jhonyy9):

yes the first triplet is 2,2,5

OpenStudy (anonymous):

asnaseer you are in the right track :)

OpenStudy (asnaseer):

I recall some property of primes that states that \(2^p-1\) where \(p\) is also prime will generate primes. is my memory correct?

OpenStudy (asnaseer):

although I'm not sure if that will help here...

OpenStudy (anonymous):

Not really but that has got some value and lead to some interesting observation check this out http://en.wikipedia.org/wiki/Mersenne_prime

OpenStudy (anonymous):

Will this help us assure that we have them all? No others as the prime numbers get larger?

OpenStudy (asnaseer):

(4,2,17) is another solution

OpenStudy (anonymous):

4 is prime ?

OpenStudy (asnaseer):

d'oh!

OpenStudy (anonymous):

Need a hint Asnaseer?

OpenStudy (asnaseer):

yes plz

OpenStudy (anonymous):

There is only solution, but how could we conclusively show that?

OpenStudy (asnaseer):

wah! - only one solution!

OpenStudy (anonymous):

No hints please. :P

OpenStudy (asnaseer):

if z is prime, then z-1 must be a composite number

OpenStudy (asnaseer):

so we get x^y=composite number

OpenStudy (anonymous):

Yes, you are close.

OpenStudy (asnaseer):

so x*x*x*...=composite number

OpenStudy (anonymous):

We know that powers of even are even numbers, and powers of odd numbers are odd numbers. The formula then suggests that if x is odd then z has to be even and vice versa.

OpenStudy (asnaseer):

which means it has two or more factors

OpenStudy (anonymous):

But we only have one even prime which is 2.

OpenStudy (anonymous):

So, if z=2, then we have \[x^y=1\] This last equation obviously has no solutions where x and y are primes.

OpenStudy (anonymous):

We're left with the case where x=2: \[2^y=z-1\]

OpenStudy (anonymous):

and soooooooo... what's the coup de grace ?

OpenStudy (anonymous):

Give me a minute.

OpenStudy (asnaseer):

any composite number can be written as :\[c=p_1*p_2*p_3*...\]where \(p_i\) are primes

OpenStudy (asnaseer):

sorry - forgot to add the powers to each prime

OpenStudy (anonymous):

The only prime p where 2^p+1 is prime is 2. We need to prove that.

OpenStudy (asnaseer):

\[c=p_1^a*p_2^b*p_3^c*...\]

OpenStudy (asnaseer):

so we get :\[x^y=p_1^a*p_2^b*p_3^c*...\]

OpenStudy (asnaseer):

and this can only be valid if:\[x^y=p_1^a\]

OpenStudy (asnaseer):

therefore no more solutions

OpenStudy (asnaseer):

not rigorous - but I think its along the right lines?

OpenStudy (anonymous):

how could you conclusively say that \( x^y=p_1^a*p_2^b*p_3^c*... \) is only valid when \( x^y=p_1^a \) ?

OpenStudy (asnaseer):

sorry I meant to say:\[x^y=p_i^j\]since x is prime. therefore it cannot be equal to the product of multiple primes.

OpenStudy (anonymous):

But how does that prove that there are no more solutions?

OpenStudy (asnaseer):

@FoolForMath: if this is a "quite easy" problem in your books then you are most definitely operating at a much higher plateau than I could ever achieve :-)

jhonyy9 (jhonyy9):

- so i think the only one solution is 2,2,5

OpenStudy (anonymous):

lol, this is really easy, I don't believe that I am any different from any of you guys :)

OpenStudy (asnaseer):

yes @jhonny9 - but how do we prove this?

OpenStudy (asnaseer):

z is prime, therefore z-1 is a composite number, therefore we can write x^y as:\[x^y=p_1^a*p_2^b*...\]but this can only be true for some \(p_i\).

OpenStudy (asnaseer):

there is some /insight/ that we are overlooking...

jhonyy9 (jhonyy9):

so my opinion is to prove this so we need to prove that ((2n+1) on exponent (2n+1) ) and added to this 1 will be always or not will be never one number in the form of (2n+1)

OpenStudy (anonymous):

I showed above that all solutions has to be of the form: \[2^y=z+1, z\ne 2.\]

jhonyy9 (jhonyy9):

(2n+1) because every primes grater than 2 can be writing in this form

OpenStudy (asnaseer):

hang on...

OpenStudy (asnaseer):

since z is prime, it must be odd (as the first solution proves) therefore z-1 must be even therefore QED?

OpenStudy (anonymous):

Yeah, z+1 must be even unless z=2, which is not a solution as I've shown above.

OpenStudy (anonymous):

So x must be even, that's x must be 2.

OpenStudy (asnaseer):

as this leads to:\[x^y=2n\]which can only be true if x=2 as that is the only even prime

OpenStudy (anonymous):

Yes.

OpenStudy (asnaseer):

but that still does not prove that there is only ONE solution :-(

OpenStudy (anonymous):

Right! It brings us closer though.

OpenStudy (asnaseer):

so this leads to x MUS be 2 as stated, so we get:\[2^y=2n\]therefore n must be a power of 2, so:\[n=2^m\]so:\[2^y=2*2^m=2^{m+1}\]therefore:\[y=m+1\]but that still does not prove only ONE solution :-(

OpenStudy (asnaseer):

working back, we get:\[z-1=2^{m+1}\]I don't know if this helps us?

OpenStudy (asnaseer):

I don't think we proved that 'm' must be prime?

OpenStudy (anonymous):

It's not, sorry!

OpenStudy (anonymous):

You guys must be thinking that I am evil :P

OpenStudy (asnaseer):

:-) - no - just smart :-)

OpenStudy (asnaseer):

ok, so m=1 gives y=2 and ALL other m's must be even for y to be a prime

OpenStudy (asnaseer):

therefore m=2k

OpenStudy (asnaseer):

therefore:\[z-1=2^{2k+1}\]and\[y=2k+1\]

OpenStudy (asnaseer):

so we need to prove that:\[z=1+2^{2k+1}\]is never prime

OpenStudy (asnaseer):

I feel like we're going in circles :-)

OpenStudy (anonymous):

Another idea, let m=z+1, then \[y=\log_{2} m\] How many m's are there where \(\log_{2} m \) is prime?

OpenStudy (anonymous):

And m-1 is prime too.

OpenStudy (anonymous):

So, m has to be in the form of \(2^n\), \(n\in Z\) in order for \(\log_{2}m\) to be integer.

OpenStudy (anonymous):

Oh sorry, m=z-1.

OpenStudy (anonymous):

Which means m+1 has to be prime.

OpenStudy (asnaseer):

@FoolForMath - are we anywhere close to the solution?

OpenStudy (anonymous):

Back again to the point where I need to show that only one prime p exists where \(2^p+1\) is prime. Although it doesn't seem that difficult to prove, I can't figure it out.

OpenStudy (anonymous):

That prime is obviously p=2.

OpenStudy (anonymous):

I have to go now, I'll be back later.

OpenStudy (asnaseer):

I /think/ I've got it. we end up with:\[z-1=2^{m+1}\]now, since z is prime, z-1 must be composite, so we can write this as:\[p_1^a*p_2^b*...=2^{m+1}\]and this can ONLY be true for one prime, namely 2.

OpenStudy (anonymous):

Probably, but you are missing a key piece of information and that why you are going round and round.

OpenStudy (anonymous):

Asnaseer, do you know that there any prime number \(\ge 5\) can be written in the form of \( 6k\pm1 \)

OpenStudy (asnaseer):

no I did not

OpenStudy (anonymous):

now you know :D :)

OpenStudy (asnaseer):

so, using that newly acquired knowledge, I get: \[2^y+1=z=6k+1\text{ OR }6k-1\]therefore:\[2^y=6k\text{ OR }6k-2\]1st solution cannot be valid, therefore:\[2^y=6k-2=2(3k-1)\]therefore:\[2^{y-1}=3k-1\]

OpenStudy (asnaseer):

but how do we prove this?

OpenStudy (anonymous):

Lets do it like this \( 2^y=6k-2\Rightarrow y=\log_2 2(3k-1) \) so, for y being prime the only possible value is when k=1. and 3k-1 is never a power of 4, Hence QED ;)

OpenStudy (asnaseer):

hmmm - I need to learn some more math before I can assert such things... :-) does this type of thing come under "Number Theory"? where would be a good place to learn?

OpenStudy (anonymous):

Yes this is number theory and University probably ? :)

OpenStudy (asnaseer):

:-) - I was just wondering if you knew any good online resources for it

OpenStudy (asnaseer):

if not - I can google for it myself.

OpenStudy (asnaseer):

ok, after spending a little more time on this I think I have an alternative proof. proving x must be 2 is straight forward and can be seen above. this leaves us with:\[2^y+1=z\]for which we see one solution as y=2, z=5 all other solutions to this MUST involve an ODD value for y (since 2 is the ONLY even prime) so let y=2n+1 to get:\[2^{2n+1}+1=z\]after trying a few values for n, I noticed that this always seems to be a multiple of 3, so I tried to prove this using induction as follows: we want to prove \(2^{2n+1}+1=3m\). We can show this is true for n=1:\[2^{2+1}+1=2^3+1=9=3*3\] lets assume its true for some n=k, so:\[2^{2k+1}+1=3m\] now lets try to prove it for n=k+1:\[2^{2(k+1)+1}+1=2^{2k+1+2}+1=4*2^{2k+1}+1\]\[=4*(3m-1)+1=12m-4+1=12m-3=3(4m-1)\]thus we have proved that \(2^{2n+1}+1\) is always a multiple of 3. therefore we can write:\[2^{2n+1}+1=z=3m\]but z is a prime number, so it cannot have factors other than 1 and itself. the only way this can be true is if m=1 giving z=3 which is not a solution to the original equation. therefore the only valid solution is x=2, y=2, z=5.

OpenStudy (asnaseer):

is my new solution sound and complete?

OpenStudy (anonymous):

and sorry, I am not aware of any online resources for number theory, if you come across one, please let me know. Thanks.

OpenStudy (anonymous):

and your new solution seems correct :) Magnum opus!